# Superlinear Convergence of an Interior Point Algorithm on Linear Semi-definite Feasibility Problems with Application to Linear Matrix Inequalities

### v3

@@ Appendix A Appendix
$\bar{\mathcal{A}}_{1}=\left(\begin{array}[]{l}{\rm{svec}}\left(\begin{array}[]{ccc}((\mathcal{A}_{1})_{1})_{11}&((\mathcal{A}_{1})_{1})_{12}&((\mathcal{A}_{1})_{1})_{13}\\ ((\mathcal{A}_{1})_{1})_{12}^{T}&((\mathcal{A}_{1})_{1})_{22}&((\mathcal{A}_{1})_{1})_{23}\\ ((\mathcal{A}_{1})_{1})_{13}^{T}&((\mathcal{A}_{1})_{1})_{23}^{T}&((\mathcal{A}_{1})_{1})_{33}\end{array}\right)\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}((\mathcal{A}_{1})_{j_{1}})_{11}&((\mathcal{A}_{1})_{j_{1}})_{12}&((\mathcal{A}_{1})_{j_{1}})_{13}\\ ((\mathcal{A}_{1})_{j_{1}})_{12}^{T}&((\mathcal{A}_{1})_{j_{1}})_{22}&((\mathcal{A}_{1})_{j_{1}})_{23}\\ ((\mathcal{A}_{1})_{j_{1}})_{13}^{T}&((\mathcal{A}_{1})_{j_{1}})_{23}^{T}&((\mathcal{A}_{1})_{j_{1}})_{33}\end{array}\right)\\ {\rm{svec}}\left(\begin{array}[]{ccc}0&((\mathcal{A}_{1})_{j_{1}+1})_{12}&0\\ ((\mathcal{A}_{1})_{j_{1}+1})_{12}^{T}&((\mathcal{A}_{1})_{j_{1}+1})_{22}&((\mathcal{A}_{1})_{j_{1}+1})_{23}\\ 0&((\mathcal{A}_{1})_{j_{1}+1})_{23}^{T}&0\end{array}\right)\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}0&((\mathcal{A}_{1})_{j_{1}+j_{2}})_{12}&0\\ ((\mathcal{A}_{1})_{j_{1}+j_{2}})_{12}^{T}&((\mathcal{A}_{1})_{j_{1}+j_{2}})_{22}&((\mathcal{A}_{1})_{j_{1}+j_{2}})_{23}\\ 0&((\mathcal{A}_{1})_{j_{1}+j_{2}})_{23}^{T}&0\end{array}\right)\\ {\rm{svec}}\left(\begin{array}[]{ccc}0&0&0\\ 0&((\mathcal{A}_{1})_{j_{1}+j_{2}+1})_{22}&0\\ 0&0&0\end{array}\right)\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}0&0&0\\ 0&((\mathcal{A}_{1})_{m_{1}})_{22}&0\\ 0&0&0\end{array}\right)\end{array}\right)$
@@
¯$\bar{\mathcal{B}}_{1}=\left(\begin{array}[]{l}{\rm{svec}}\left(\begin{array}[]{ccc}((\mathcal{B}_{1})_{1})_{11}&((\mathcal{B}_{1})_{1})_{12}&((\mathcal{B}_{1})_{1})_{13}\\ ((\mathcal{B}_{1})_{1})_{12}^{T}&((\mathcal{B}_{1})_{1})_{22}&((\mathcal{B}_{1})_{1})_{23}\\ ((\mathcal{B}_{1})_{1})_{13}^{T}&((\mathcal{B}_{1})_{1})_{23}^{T}&((\mathcal{B}_{1})_{1})_{33}\end{array}\right)\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}((\mathcal{B}_{1})_{k_{1}})_{11}&((\mathcal{B}_{1})_{k_{1}})_{12}&((\mathcal{B}_{1})_{k_{1}})_{13}\\ ((\mathcal{B}_{1})_{k_{1}})_{12}^{T}&((\mathcal{B}_{1})_{k_{1}})_{22}&((\mathcal{B}_{1})_{k_{1}})_{23}\\ ((\mathcal{B}_{1})_{k_{1}})_{13}^{T}&((\mathcal{B}_{1})_{k_{1}})_{23}^{T}&((\mathcal{B}_{1})_{k_{1}})_{33}\end{array}\right)\\ {\rm{svec}}\left(\begin{array}[]{ccc}((\mathcal{B}_{1})_{k_{1}+1})_{11}&((\mathcal{B}_{1})_{k_{1}+1})_{12}&((\mathcal{B}_{1})_{k_{1}+1})_{13}\\ ((\mathcal{B}_{1})_{k_{1}+1})_{12}^{T}&0&((\mathcal{B}_{1})_{k_{1}+1})_{23}\\ ((\mathcal{B}_{1})_{k_{1}+1})_{13}^{T}&((\mathcal{B}_{1})_{k_{1}+1})_{23}^{T}&((\mathcal{B}_{1})_{k_{1}+1})_{33}\end{array}\right)\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}((\mathcal{B}_{1})_{k_{1}+k_{2}})_{11}&((\mathcal{B}_{1})_{k_{1}+k_{2}})_{12}&((\mathcal{B}_{1})_{k_{1}+k_{2}})_{13}\\ ((\mathcal{B}_{1})_{k_{1}+k_{2}})_{12}^{T}&0&((\mathcal{B}_{1})_{k_{1}+k_{2}})_{23}\\ ((\mathcal{B}_{1})_{k_{1}+k_{2}})_{13}^{T}&((\mathcal{B}_{1})_{k_{1}+k_{2}})_{23}^{T}&((\mathcal{B}_{1})_{k_{1}+k_{2}})_{33}\end{array}\right)\\ {\rm{svec}}\left(\begin{array}[]{ccc}((\mathcal{B}_{1})_{k_{1}+k_{2}+1})_{11}&0&((\mathcal{B}_{1})_{k_{1}+k_{2}+1})_{13}\\ 0&0&0\\ ((\mathcal{B}_{1})_{k_{1}+k_{2}+1})_{13}^{T}&0&((\mathcal{B}_{1})_{k_{1}+k_{2}+1})_{33}\end{array}\right)\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}((\mathcal{B}_{1})_{\tilde{n}_{1}-m_{1}})_{11}&0&((\mathcal{B}_{1})_{\tilde{n}_{1}-m_{1}})_{13}\\ 0&0&0\\ ((\mathcal{B}_{1})_{\tilde{n}_{1}-m_{1}})_{13}^{T}&0&((\mathcal{B}_{1})_{\tilde{n}_{1}-m_{1}})_{33}\end{array}\right)\end{array}\right).$
Details on the above reformulation of $\mathcal{A}_{1}$ and $\mathcal{B}_{1}$ can be found in [26, 27, 28].
@@
In the context of our SDLCP representation of the homogeneous feasibility model of the given LSDFP, we choose $B_{1}$ in (55) such that $(B_{1})_{22}\not=0$. In this way, when we perform block Gaussian eliminations and row operations on $\mathcal{B}_{1}$, its final position with respect to other rows in ¯$\bar{\mathcal{B}}_{1}$ remains unchanged, that is, ${\rm{svec}}\left(\begin{array}[]{cc}B_{1}&0\\ 0&d_{1}\end{array}\right)^{T}$ remains in the first row of ¯$\bar{\mathcal{B}}_{1}$ or the $(m_{1}+1)^{th}$ row of $\mathcal{B}_{1}$. We see that (155) then holds when $Y^{1}_{0}$ is from a strictly feasible solution to the dual SDP (8) with $C=0$.

### v4

@@ Appendix A Appendix
${{\color[rgb]{0,0,0}\bar{\mathcal{A}}}}=\left(\begin{array}[]{l}{\rm{svec}}\left(\begin{array}[]{ccc}(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{1})_{11}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{1})_{12}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{1})_{13}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{1})_{12}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{1})_{22}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{1})_{23}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{1})_{13}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{1})_{23}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{1})_{33}\end{array}\right)^{T}\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}})_{11}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}})_{12}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}})_{13}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}})_{12}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}})_{22}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}})_{23}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}})_{13}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}})_{23}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}})_{33}\end{array}\right)^{T}\\ {\rm{svec}}\left(\begin{array}[]{ccc}0&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+1})_{12}&0\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+1})_{12}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+1})_{22}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+1})_{23}\\ 0&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+1})_{23}^{T}&0\end{array}\right)^{T}\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}0&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+j_{2}})_{12}&0\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+j_{2}})_{12}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+j_{2}})_{22}&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+j_{2}})_{23}\\ 0&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+j_{2}})_{23}^{T}&0\end{array}\right)^{T}\\ {\rm{svec}}\left(\begin{array}[]{ccc}0&0&0\\ 0&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+j_{2}+1})_{22}&0\\ 0&0&0\end{array}\right)^{T}\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}0&0&0\\ 0&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{m_{1}})_{22}&0\\ 0&0&0\end{array}\right)^{T}\end{array}\right)$
@@
¯${{\color[rgb]{0,0,0}\bar{\mathcal{B}}}}=\left(\begin{array}[]{l}{\rm{svec}}\left(\begin{array}[]{ccc}(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{1})_{11}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{1})_{12}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{1})_{13}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{1})_{12}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{1})_{22}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{1})_{23}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{1})_{13}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{1})_{23}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{1})_{33}\end{array}\right)^{T}\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}})_{11}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}})_{12}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}})_{13}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}})_{12}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}})_{22}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}})_{23}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}})_{13}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}})_{23}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}})_{33}\end{array}\right)^{T}\\ {\rm{svec}}\left(\begin{array}[]{ccc}(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+1})_{11}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+1})_{12}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+1})_{13}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+1})_{12}^{T}&0&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+1})_{23}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+1})_{13}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+1})_{23}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+1})_{33}\end{array}\right)^{T}\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}})_{11}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}})_{12}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}})_{13}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}})_{12}^{T}&0&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}})_{23}\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}})_{13}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}})_{23}^{T}&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}})_{33}\end{array}\right)^{T}\\ {\rm{svec}}\left(\begin{array}[]{ccc}(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}+1})_{11}&0&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}+1})_{13}\\ 0&0&0\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}+1})_{13}^{T}&0&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}+1})_{33}\end{array}\right)^{T}\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{\tilde{n}_{1}-m_{1}})_{11}&0&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{\tilde{n}_{1}-m_{1}})_{13}\\ 0&0&0\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{\tilde{n}_{1}-m_{1}})_{13}^{T}&0&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{\tilde{n}_{1}-m_{1}})_{33}\end{array}\right)^{T}\end{array}\right).$
Details on the above reformulation of ${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}$ and $^{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}$ can be found in [26, 27, 28].
@@
In the context of our SDLCP representation, sSDLCP, of the homogeneous feasibility model of the sLSDFP, where we have $m_{1}=m+n,n_{1}=n+1$, and $\hat{\mathcal{A}},\hat{\mathcal{B}}$ are given by (73), (78) respectively, we choose $B_{1}$ such that $(B_{1})_{22}\not=0$. In this way, when we perform block Gaussian eliminations and row operations on $^{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}$, its final position with respect to other rows in ${{\color[rgb]{0,0,0}\bar{\mathcal{B}}}}$ remains unchanged, that is, ${\rm{svec}}\left(\begin{array}[]{cc}B_{1}&0\\ 0&d_{1}\end{array}\right)^{T}$ remains in the first row of ${{\color[rgb]{0,0,0}\bar{\mathcal{B}}}}$ or the $(m_{1}+1)^{th}$ row of $^{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}$. We see that (148) (which is (52) of Theorem 5.1 in [26]) then holds when ${{\color[rgb]{0,0,0}\hat{Y}}}_{0}$ is from a strictly feasible solution to the sDSDP.